OpenStudy (anonymous):

What is 15 over 8 as a mixed number

OpenStudy (anonymous):

1 and 7/8's

OpenStudy (unklerhaukus):

\[\frac{a+b}{a}=\frac{a}{a}+\frac{b}{a}=1+\frac ba\]

OpenStudy (anonymous):

or 15/8

OpenStudy (anonymous):

1 7/8

OpenStudy (anonymous):

simplified yes 1 7/8

OpenStudy (anonymous):

yep